# Spin Polarization of $\boldsymbol{\Lambda}$ hyperons from Dissipative Spin Hydrodynamics

### v1

@@ Abstract
Summary.– In this work, we solved the equations of relativistic dissipative spin hydrodynamics on a (3+1)-dimensional relativistic hydrodynamic background for the first time. The evolution equations for spin-$\nicefrac{{1}}{{2}}$ particles, derived in Ref. [53], using quantum kinetic theory with non-local collisions, were employed in our study. We assumed zero baryon diffusion in both the hydrodynamic background and the spin evolution equations, and considered the particles to have a constant mass of 300 MeV. While the first of these assumptions can be lifted by including baryon diffusion, weakening the second one requires, e.g., considering temperature-dependent masses, which would introduce additional source terms in the spin evolution equations [75]. We leave the study of such effects to future work.
@@
Our results show that the relevant spin degrees of freedom ($\omega_{0}^{\mu}$, $\kappa_{0}^{\mu}$ and $\mathfrak{t}^{\mu\nu}$), initially set to zero, are generated through the exchange of orbital and spin angular momentum. The initial orbital angular momentum originates from the background hydrodynamic initial condition, and the spin degrees of freedom subsequently evolve via spin-orbit interactions among particles. The nature of the microscopic interaction is encoded in the transport coefficients of spin hydrodynamics. We have considered three types of interactions between the particles, which primarily affect the relaxation times $\tau_{\omega}$, $\tau_{\kappa}$, $\tau_{\mathfrak{t}}$, as well as the transport coefficient $\mathfrak{d}$. All other transport coefficients remain independent of interactions when expressed in units of appropriate combinations of the respective relaxation times and pressure. We apply this framework to study spin dynamics in the hot and dense fireball formed in relativistic HIC. For $\sqrt{s_{NN}}=200$ GeV, our results indicate that the spin potential relaxes to the thermal vorticity during the fireball’s lifetime for the interactions considered in this work. A non-zero spin-shear stress ($\mathfrak{t}^{\mu\nu}$) accelerates this relaxation while also reducing the peak value of the spin potential, acting as a dissipative effect. The framework is then used to compute the spin polarization of $\Lambda$ hyperons. Our predictions show reasonable agreement with experimental data for both global and local polarization for the $I_{SP}$ interaction scenario. Notably, our approach not only reproduces the correct sign of the longitudinal polarization but also demonstrates its sensitivity to microscopic details, while the global polarization remains only slightly affected by them. The only free parameter of our approach is the particle mass, highlighting the predictive power of the theory. Our analysis suggests that the inclusion of low temperature fluid cells delays spin relaxation. This effect becomes particularly significant at lower collision energies, where a baryon-rich fireball is produced at low temperatures, leading to longer spin-relaxation times. Consequently, the spin may not fully thermalize within the fireball’s lifetime, making our framework especially relevant to such conditions.
@@ Appendix A Background Hydrodynamics
$\dot{\pi}^{\langle\alpha\beta\rangle}$ $=\frac{\pi_{\rm NS}^{\alpha\beta}-\pi^{\alpha\beta}}{\tau_{\pi}}-\frac{\delta_{\pi\pi}}{\tau_{\pi}}\pi^{\alpha\beta}\theta+\frac{\lambda_{\pi\Pi}}{\tau_{\pi}}\Pi\,\sigma^{\alpha\beta}-\frac{\tau_{\pi\pi}}{\tau_{\pi}}\pi_{\gamma}^{\langle\alpha}\sigma^{\beta\rangle\gamma}+\frac{\varphi_{7}}{\tau_{\pi}}\pi_{\gamma}^{\langle\alpha}\pi^{\beta\rangle\gamma}\;.$ (11)
@@
$\tau_{\pi}=\frac{5\eta}{\varepsilon+P}\;,\quad\tau_{\Pi}=\frac{\zeta}{15\left(\frac{1}{3}-c_{s}^{2}\right)^{2}(\varepsilon+P)}\;,$ (14)
$\frac{\delta_{\Pi\Pi}}{\tau_{\Pi}}=\frac{2}{3}\;,\quad\frac{\lambda_{\Pi\pi}}{\tau_{\Pi}}=\frac{8}{5}\left(\frac{1}{3}-c_{s}^{2}\right)\;,\quad\frac{\delta_{\pi\pi}}{\tau_{\pi}}=\frac{4}{3}\;,$ (15)
@@
For the initial energy-momentum tensor, we use the model of Refs [81, 82], which is based on the Glauber collision geometry and local energy-momentum conservation, with the only difference being that we compute the thickness functions and wounded nucleon densities using the optical limit of the Glauber model (cf. Appendix A of Ref. [54] for details). The bulk viscous pressure ($\Pi$) and the shear-stress tensor ($\pi^{\mu\nu}$) are initialized to zero. In this work, we focus on the Au+Au collisions at the top RHIC energy of $\sqrt{s_{\rm NN}}=200$ GeV. The impact parameter $b$ corresponding to various centralities are taken from Ref. [54]. The initial time for hydrodynamic evolution is fixed at $\tau_{0}=$1 fm.
A description of observables, such as momentum spectra and flow coefficients, does not require the evolution of spin equations, as these are decoupled from the background. The initial condition model parameters and the switching energy density for the particlization surface, where fluid cells are converted to hadrons, are estimated by fitting the model predictions to experimental data. This calibration has already been performed in Ref. [54] for $\sqrt{s_{\rm NN}}=200$ GeV, and we use the same set of parameters in this work. The particlization surface is obtained using the CORNELIUS code [83]. At the particlization surface, we use a hadron sampler [84, 85, 86] to generate particles from fluid elements. The resulting particle set serves as input to the SMASH transport model [87, 88] for subsequent hadron interactions and decays. To validate the background hydrodynamic code, we re-run our simulations and reproduce the results of Ref. [54] for three centrality classes, as shown in Fig. 3.
@@ Appendix B Spin relaxation times and transport coefficients
The spin relaxation times ($\tau_{\omega}$, $\tau_{\kappa}$, $\tau_{\mathfrak{t}}$) and the first-order transport coefficient $\mathfrak{d}$, which are used in this work, are plotted as functions of $z$ in Fig. 4 for three different interactions. All other transport coefficients in Eqs. (2a)-(2c) are discussed in Ref. [53].
@@ Appendix C Definition of spatial average
Figure 6: Time evolution of the energy-weighted space-averaged $xz$- component of the spin potential for scalar interaction and for $\varepsilon_{\text{cut}}=0.1$ GeV/fm3.
@@ Appendix D Numerical details
We denote the value of $\psi$ at position $(x_{i},y_{j},{\eta_{s}}{}_{,k})$ and at $n^{\text{th}}$ time step by $\psi_{ijk}^{n}$, i.e., $\psi_{ijk}^{n}=\psi(t_{n},x_{i},y_{j},{\eta_{s}}{}_{,k})$. We first solve Eq. (19) according to the following prescription
@@
In order to cast Eqs. (2a)-(2c) in the form of Eq. (18), we follow the notation of Ref. [61] and define a scaled four-vector, $\tilde{A}^{\mu}$, using a given four-vector $A^{\mu}$ as follows:
@@
$\begin{array}[]{cccl}\widetilde{S}^{\mu\nu}&=&S^{\mu\nu}&\text{ for }\mu,\nu\neq{\eta_{s}}\;,\\ \widetilde{S}^{\mu\nu}&=&\tau S^{\mu\nu}&\text{ if any one of }\ \mu,\nu\text{is }{\eta_{s}}\;,\\ \widetilde{S}^{{\eta_{s}}{\eta_{s}}}&=&\tau^{2}S^{{\eta_{s}}{\eta_{s}}}\;.&\end{array}$
@@
The equations that we solve numerically are obtained from Eqs. (2a)-(2c) after re-writing in the above notation and are given by:
@@
Figure 7: Comparison of results from dissipative spin hydrodynamics for $I_{SP}$ interaction scenario with BBP prescription for Au+Au collisions at $\sqrt{s_{NN}}=200$ GeV. The simulations are performed with fixed impact parameter, $b=8.4$ fm, corresponding to 20-50% centrality. The top panel shows the (left) pseudorapidity and (right) transverse momentum dependence of the $y$-component of $\Lambda$ polarization. The bottom panel shows the (left) azimuthal and (right) transverse momentum dependence of the $z$-component of $\Lambda$ polarization. Experimental data are taken from Refs. [2, 72] with the updated decay parameter ($\alpha_{\Lambda}=0.732$).
@@ Appendix E Comparison to quantum-statistical approach
We compare our results on $\Lambda$-polarization, calculated using Eq. (3), with the results obtained using the prescription from Refs. [11, 13], which is given by
@@
Figure 8: Individual contributions from $\omega_{0}^{\mu}$, $\kappa_{0}^{\mu}$ and $\mathfrak{t}^{\mu\nu}$ to the (left) $y$-component and (right) $z$-component of $\Lambda$ polarization for three different interactions. Note that $P^{z}_{\omega}$ is very small and appears as zero on the chosen scale of the plot, which is kept consistent to allow a visual comparison of the relative individual contributions.
Without any approximation, the longitudinal polarization due to combined effects of thermal vorticity and thermal shear is negative, as shown by green dash-dotted lines in lower panel of Fig. 7. With the isothermal approximation [13], the surface $\Sigma$ is assumed to be at constant temperature (which is a good approximation at high energies), and factors of temperature are pulled out of the integral in the density operator rather than expanded in gradients. Under this approximation, defining $\Xi^{\mu\nu}=\frac{1}{2}\partial^{(\mu}u^{\nu)}$ and $\omega^{\mu\nu}=\frac{1}{2}\partial^{[\mu}u^{\nu]}$, the expressions (28) become
@@
The combined effects of thermal vorticity and thermal shear in the isothermal approximation gives the correct sign of longitudinal polarization, as shown by blue dashed lines in lower panel of Fig. 7. It is worth noting that the isothermal approximation employed in Ref. [13], if it were applied to the spin-hydrodynamic equations (2a)–(2c), involves neglecting the temperature gradient term, which would result in an asymptotic value of $\kappa_{0,\text{NS,iso}}^{\mu}=u_{\nu}\varpi^{\nu\mu}\approx-\dot{u}^{\mu}/(2T)$ – half of the value obtained without the approximation.
@@ Appendix F Additional figures
In the left panel of Fig. 8, we show the individual contributions from $\omega^{\mu}_{0}$, $\kappa^{\mu}_{0}$ and $\mathfrak{t}^{\mu\nu}$ to the transverse momentum ($p_{T}$) dependence of the $y$-component of $\Lambda$ polarization for three different interactions, and in the right panel of Fig. 8, we show the individual contributions to the azimuthal angle dependence of the $z$-component of $\Lambda$ polarization. The net spin polarization is obtained by summing the individual contributions.
In the main text, we fixed the impact parameter of Au+Au collisions to 8.4 fm, corresponding to the mean impact parameter in the 20–50% centrality class for Au+Au collision at $\sqrt{s_{NN}}=200$ GeV. In Fig. 9, we show the centrality dependence of the $y$- and $z$-components of $\Lambda$-spin polarization. Results from dissipative spin hydrodynamics for the $I_{SP}$ interaction scenario agree well with experimental data.
In Fig. 10, we show the transverse momentum dependence of the $y$-component of polarization and the azimuthal angle dependence of the $z$-component of polarization for three different masses. The curves for masses of 300 and 500 MeV are nearly identical. In contrast, the sign of the longitudinal polarization is negative for $m=100$ MeV. These results can be understood by examining the $\mathfrak{d}$ vs. $z$ plot in Fig. 4. For $m=100$ MeV, $\langle z\rangle$ of the fireball varies between 0.4 and 0.62, with a smaller $\mathfrak{d}$, leading to reduced dissipative effects. For $m=500$ MeV, $\langle z\rangle$ varies between 2.0 and 3.1, resulting in a similar value of $\mathfrak{d}$, compared to the case of $m=300$ MeV.
Figure 10: For the $I_{SP}$ interaction scenario, the left panel shows the transverse momentum dependence of the $y$-component of polarization, while the right panel shows the azimuthal angle dependence of $z$-component for three different masses.
Experimental data are taken from Refs. [2, 72] with the updated decay parameter ($\alpha_{\Lambda}=0.732$).
In Fig. 1, we demonstrated that the spin potential approaches its asymptotic value for all three interactions (left panel). However, for the spin-shear stress shown (right panel), we did not show the asymptotic values, as they differ for the three interactions due to the transport coefficient $\mathfrak{d}$. In Fig. 11, we demonstrate for the $I_{SP}$ scenario that the spin-shear stress does approach its asymptotic value $\frac{\mathfrak{d}}{T}\sigma^{\mu\nu}$.

### v2

@@ V Summary
In this work, we solved the equations of relativistic dissipative spin hydrodynamics on a (3+1)-dimensional relativistic hydrodynamic background. The evolution equations for spin-$\nicefrac{{1}}{{2}}$ particles, derived in Ref. [55], using quantum kinetic theory with nonlocal collisions, were employed in our study. We assumed zero baryon diffusion in both the hydrodynamic background and the spin-evolution equations, and considered the particles to have a constant mass of 300 MeV. While the first of these assumptions can be lifted by including baryon diffusion, weakening the second one requires, e.g., considering temperature-dependent masses, which would introduce additional source terms in the spin evolution equations [79]. We leave the study of such effects to future work.
@@
Our results show that the relevant spin degrees of freedom ($\omega_{0}^{\mu}$, $\kappa_{0}^{\mu}$ and $\mathfrak{t}^{\mu\nu}$), initially set to zero, are generated through the exchange of orbital and spin angular momentum. The initial orbital angular momentum originates from the background hydrodynamic initial condition, and the spin degrees of freedom subsequently evolve via spin-orbit interactions among particles. The nature of the microscopic interaction is encoded in the transport coefficients of spin hydrodynamics. We have considered three types of interactions between the particles, which primarily affect the relaxation times $\tau_{\omega}$, $\tau_{\kappa}$, $\tau_{\mathfrak{t}}$, as well as the transport coefficient $\mathfrak{d}$. All other transport coefficients remain independent of interactions when expressed in units of appropriate combinations of the respective relaxation times and pressure. We apply this framework to study spin dynamics in the hot and dense fireball formed in relativistic HICs. For $\sqrt{s_{NN}}=200$ GeV, our results indicate that the spin potential relaxes to the thermal vorticity during the fireball’s lifetime for the interactions considered in this work. A nonzero spin-shear stress ($\mathfrak{t}^{\mu\nu}$) accelerates this relaxation while also reducing the peak value of the spin potential, acting as a dissipative effect. The framework is then used to compute the spin polarization of $\Lambda$ hyperons. Our predictions show reasonable agreement with experimental data for both global and local polarization for the $I_{SP}$ interaction scenario. Notably, our approach not only reproduces the correct sign of the longitudinal polarization but also demonstrates its sensitivity to microscopic details, while the global polarization remains only slightly affected by them. The only free parameter in the equations of spin hydrodynamics is the particle mass, highlighting the predictive power of the theory. Our analysis suggests that the inclusion of low temperature fluid cells delays spin relaxation. This effect becomes particularly significant at lower collision energies, where a baryon-rich fireball is produced at low temperatures, leading to longer spin-relaxation times. Consequently, the spin may not fully thermalize within the fireball’s lifetime, making our framework especially relevant to such conditions.
@@ Appendix A Background Hydrodynamics
$\dot{\pi}^{\langle\alpha\beta\rangle}$ $=\frac{\pi_{\rm NS}^{\alpha\beta}-\pi^{\alpha\beta}}{\tau_{\pi}}-\frac{\delta_{\pi\pi}}{\tau_{\pi}}\pi^{\alpha\beta}\theta+\frac{\lambda_{\pi\Pi}}{\tau_{\pi}}\Pi\,\sigma^{\alpha\beta}$
$\quad-\frac{\tau_{\pi\pi}}{\tau_{\pi}}\pi_{\gamma}^{\langle\alpha}\sigma^{\beta\rangle\gamma}+\frac{\varphi_{7}}{\tau_{\pi}}\pi_{\gamma}^{\langle\alpha}\pi^{\beta\rangle\gamma}\;.$ (11)
@@
$\tau_{\pi}=\frac{5\eta}{\varepsilon+P}\;,\quad\tau_{\Pi}=\frac{\zeta}{15\left(\frac{1}{3}-c_{s}^{2}\right)^{2}(\varepsilon+P)}\;,$
$\frac{\delta_{\Pi\Pi}}{\tau_{\Pi}}=\frac{2}{3}\;,\quad\frac{\lambda_{\Pi\pi}}{\tau_{\Pi}}=\frac{8}{5}\left(\frac{1}{3}-c_{s}^{2}\right)\;,\quad\frac{\delta_{\pi\pi}}{\tau_{\pi}}=\frac{4}{3}\;,$
@@
For the initial energy-momentum tensor, we use the model of Refs [88, 89], which is based on the Glauber collision geometry and local energy-momentum conservation, with the only difference being that we compute the thickness functions and wounded-nucleon densities using the optical limit of the Glauber model (cf. Appendix A of Ref. [59] for details). The bulk viscous pressure ($\Pi$) and the shear-stress tensor ($\pi^{\mu\nu}$) are initialized to zero. In this work, we focus on the Au+Au collisions at the top RHIC energy of $\sqrt{s_{\rm NN}}=200$ GeV. The impact parameter $b$ corresponding to various centralities are taken from Ref. [59]. The initial time for hydrodynamic evolution is fixed at $\tau_{0}=$1 fm.
A description of observables, such as momentum spectra and flow coefficients, does not require the evolution of spin equations, as these are decoupled from the background. The initial condition model parameters and the switching energy density for the particlization surface, where fluid cells are converted to hadrons, are estimated by fitting the model predictions to experimental data. This calibration has already been performed in Ref. [59] for $\sqrt{s_{\rm NN}}=200$ GeV, and we use the same set of parameters in this work. The particlization surface is obtained using the CORNELIUS code [90]. At the particlization surface, we use a hadron sampler [91, 92, 93] to generate particles from fluid elements. The resulting particle set serves as input to the SMASH transport model [94, 95] for subsequent hadron interactions and decays. To validate the background hydrodynamic code, we rerun our simulations and reproduce the results of Ref. [59] for three centrality classes, as shown in Fig. 3.
@@ Appendix B Spin relaxation times and transport coefficients
The spin relaxation times ($\tau_{\omega}$, $\tau_{\kappa}$, $\tau_{\mathfrak{t}}$) and the first-order transport coefficient $\mathfrak{d}$, which have been computed in a Mathematica notebook [55, 96, 97, 98, 99], are plotted as functions of $z=m/T$ in Fig. 4 for three different interactions. All other transport coefficients in the evolution equations for the spin degrees of freedom $(\omega_{0}^{\mu},\ \kappa_{0}^{\mu},\ \mathfrak{t}^{\mu\nu})$ are discussed in Ref. [55].
@@ Appendix C Definition of spatial average
Figure 6: Time evolution of the energy-weighted space-averaged xzxz component of the spin potential for scalar interaction and for $\varepsilon_{\text{cut}}=0.1$ GeV/fm3.
@@ Appendix D Numerical details
Figure 7: Comparison of results from dissipative spin hydrodynamics for $I_{SP}$ interaction scenario with BBP prescription for Au+Au collisions at $\sqrt{s_{NN}}=200$ GeV. The simulations are performed with fixed impact parameter, $b=8.4$ fm, corresponding to 20-50% centrality. The top panel shows the (left) pseudorapidity and (right) transverse momentum dependence of the $y$ component of $\Lambda$ polarization. The bottom panel shows the (left) azimuthal and (right) transverse momentum dependence of the $z$ component of $\Lambda$ polarization. Experimental data are taken from Refs. [2, 3] with the updated decay parameter ($\alpha_{\Lambda}=0.732$).
@@
[…]
@@
Figure 8: Individual contributions from $\omega_{0}^{\mu}$, $\kappa_{0}^{\mu}$ and $\mathfrak{t}^{\mu\nu}$ to the (left) $y$ component and (right) $z$ component of $\Lambda$ polarization for three different interactions. Note that $P^{z}_{\omega}$ is very small and appears as zero on the chosen scale of the plot, which is kept consistent to allow a visual comparison of the relative individual contributions.
In order to cast the evolution equations for the spin degrees of freedom $(\omega_{0}^{\mu},\ \kappa_{0}^{\mu},\ \mathfrak{t}^{\mu\nu})$ in the form of Eq. (16), we follow the notation of Ref. [66] and define a scaled four-vector, $\tilde{A}^{\mu}$, using a given four-vector $A^{\mu}$ as follows:
@@
[…]
@@
The equations that we solve numerically are obtained from spin evolution equations after re-writing in the above notation and are given by:
@@ Appendix E Comparison to quantum-statistical approach
We compare our results for $\Lambda$-polarization, calculated within our framework [using Eq. (3) of main text], with those obtained using the prescription from Refs. [11, 13], which is given by
@@
Without any approximation, the longitudinal polarization due to the combined effects of thermal vorticity and thermal shear is negative, as shown by green dash-dotted lines in lower panel of Fig. 7. With the isothermal approximation [13], the surface $\Sigma$ is assumed to be at constant temperature (which is a good approximation at high energies), and factors of temperature are pulled out of the integral in the density operator rather than expanded in gradients. Under this approximation, defining $\Xi^{\mu\nu}=\frac{1}{2}\partial^{(\mu}u^{\nu)}$ and $\omega^{\mu\nu}=\frac{1}{2}\partial^{[\mu}u^{\nu]}$, the expressions (26) become
@@
The combined effects of thermal vorticity and thermal shear in the isothermal approximation give the correct sign of longitudinal polarization, as shown by blue dashed lines in lower panel of Fig. 7. It is worth noting that the isothermal approximation employed in Ref. [13], if it were applied to the evolution equations for the spin degrees of freedom $(\omega_{0}^{\mu},\ \kappa_{0}^{\mu},\ \mathfrak{t}^{\mu\nu})$, involves neglecting the temperature-gradient term, which would result in an asymptotic value of $\kappa_{0,\text{NS,iso}}^{\mu}=u_{\nu}\varpi^{\nu\mu}\approx-\dot{u}^{\mu}/(2T)$ – half of the value obtained without the approximation.
@@ Appendix F Additional figures
In the left panel of Fig. 8, we show the individual contributions from $\omega^{\mu}_{0}$, $\kappa^{\mu}_{0}$ and $\mathfrak{t}^{\mu\nu}$ to the transverse momentum ($p_{T}$) dependence of the $y$ component of $\Lambda$ polarization for three different interactions, and in the right panel of Fig. 8, we show the individual contributions to the azimuthal angle dependence of the $z$ component of $\Lambda$ polarization. The net spin polarization is obtained by summing the individual contributions.
In the main text, we fixed the impact parameter of Au+Au collisions to 8.4 fm, corresponding to the mean impact parameter in the 20–50% centrality class for Au+Au collision at $\sqrt{s_{NN}}=200$ GeV. In Fig. 9, we show the centrality dependence of the $y$ and $z$ components of $\Lambda$-spin polarization. Results from dissipative spin hydrodynamics for the $I_{SP}$ interaction scenario agree well with experimental data.
In Fig. 10, we show the transverse momentum dependence of the $y$ component of polarization and the azimuthal angle dependence of the $z$ component of polarization for three different masses. The curves for masses of 300 and 500 MeV are nearly identical. In contrast, the sign of the longitudinal polarization is negative for $m=100$ MeV. These results can be understood by examining the $\mathfrak{d}$ vs. $z$ plot in Fig. 4. For $m=100$ MeV, $\langle z\rangle$ of the fireball varies between 0.4 and 0.62, with a smaller $\mathfrak{d}$, leading to reduced dissipative effects. For $m=500$ MeV, $\langle z\rangle$ varies between 2.0 and 3.1, resulting in a similar value of $\mathfrak{d}$, compared to the case of $m=300$ MeV.
Figure 10: For the $I_{SP}$ interaction scenario, the left panel shows the transverse momentum dependence of the $y$ component of polarization, while the right panel shows the azimuthal angle dependence of $z$ component for three different masses. Experimental data are taken from Refs. [2, 3] with the updated decay parameter ($\alpha_{\Lambda}=0.732$).
@@
In the main text, we demonstrated that the spin potential approaches its asymptotic value for all three interactions. However, for the spin-shear stress, we did not show the asymptotic values, as they differ for the three interactions due to the transport coefficient $\mathfrak{d}$. In Fig. 11, we demonstrate for the $I_{SP}$ scenario that the spin-shear stress does approach its asymptotic value $\frac{\mathfrak{d}}{T}\sigma^{\mu\nu}$.